What is the potential energy if a +4 nC charge moves from infinity to point A, 5 cm away from a +3 μC charge?


1
Expert's answer
2022-02-25T11:12:25-0500

Potential energy:


"U=\frac{kQq}{r}"

"U=\frac{(9×10^9\frac{Nm^2}{C^2})(+3×10^{-6}C)(+4×10^{-9}C}{(0.05m)}"